OpenStudy (anonymous):

Factorise 7x-14 What is the answer?

OpenStudy (mathmale):

@chaudhry.s: Please understand that OpenStudy is not a source of free answers. You may ask for guidance towards solving a problem, but not for the answer itself. And other users are not allowed to give out answers with no work on the part of the person asking the question.

OpenStudy (mathmale):

Note that 7x-14 is not an equation (at least, not the way you have typed it). To factor this expression, look for the factor common to 7x and 14. What is it? Factor this out.

OpenStudy (anonymous):

Thank you @engrbill @mathmale I needed a way of working it out

OpenStudy (mathmale):

@engrbill: Please note that unless chaudhry.s tells us otherwise, this is an expression, not an equation, and if it is not an equation, there is no "solution." "And other users are not allowed to give out answers with no work on the part of the person asking the question."

OpenStudy (mathmale):

Note that 7x-14 is not an equation (at least, not the way you have typed it). @chaudhry.s: I have already given you some instructions: "To factor this expression, look for the factor common to 7x and 14. What is it? Factor this out."

OpenStudy (anonymous):

@ mathmale this is also a non-ethical way to remove the comment of any user without his permission...

OpenStudy (anonymous):

Yeah @mathmale why did you remove @engrbill 's comment??

OpenStudy (mathmale):

@engrbill: before you accuse me of being unethical, please review the Code of Conduct and see what it has to say about either asking just for answers or doing all of someone else's work and presenting the answer without any invovlement on the part of the questioner. And once more: this is not an equation, but an expression, so if you treat chaudhry.s ' post as an equation, you are misleading him.

OpenStudy (mathmale):

@chaudhry.s: I have already given you some instructions: "To factor this expression, look for the factor common to 7x and 14. What is it? Factor this out."

OpenStudy (anonymous):

indeed today i am very shocked by the action of someone... at least this is an educational site,,, not a political one... and mr mathmale everyone has his opinion.... and it will be better for you if you rectify my answer, instead of deleting sir...

OpenStudy (anonymous):

@mathmale : You didn't need to delete @engrbill 's comment, unless you have his permission! Also did you call me a man?

OpenStudy (mathmale):

This is a private educational site, as you point out. As such, it has rules, contained in a Code of Conduct, and if you wish to use this site, you are bound to follow those rules. Further, as a Moderator it is my job to enforce those rules. My actions are not political. You have had fifteen minutes or more to guide chaudry.s towards finding the correct solution to his problem, but instead have chosen to argue where really you have no grounds for argument (in view of the reality of this site being a private, educational site). I have no idea what your gender is, and that's rather irrelevant, isn't it?

OpenStudy (mathmale):

either help chaudhry.s towards finding the correct answer to the problem posted, or STOP posting...unless you want to be accused of spamming.

OpenStudy (anonymous):

@ mathmale hats off for you sir... and also for your reply... just ask from the owner of this question, that when i had replied him, after the question post... and sure, i am following the rules as i am a man of rules...

OpenStudy (anonymous):

@engrbill I'm a female. :)

OpenStudy (anonymous):

by the way thanks for wasting my time... i came here to learn and to explain... not for reasoning sir....

OpenStudy (mathmale):

I see no instructions, no help from you, towards factoring the expression posted by chaudhry.s. So, @engrbill, either provide them or stop spamming this site. Note, once more, that 7x-14 is not an equation, and therefore does not have a "solution"; instead, 7x-14 has factors. Explain to chaudhry.s, please, how to factor that expression.

OpenStudy (anonymous):

@chaudhry.s mam i have no interest with your gender... and also i haven't mentioned your gender throughout the discussion....

OpenStudy (anonymous):

mathmale ok sir... sorry... i am wrong... now happy :-)

OpenStudy (mathmale):

@engrbill: That was your last chance. Since you haven't helped chaudhry.s yet, I will do it. @chaudhry.s: Your expression, 7x-14, has two terms. 7 is a factor of each of these terms, as shown here:\[7x-14=7(x) - 7(2)\] Therefore, we factor the 7 out. We get 7x-14=7( ??? )

OpenStudy (anonymous):

@ mathmale is it right one or not... if we take common the 7 from the expression... and solve this like this... 7x-14=0 7(x-2)=0 here x=2